-8x-20=25 and 1/3
x=?

Answer by algebrahouse.com(1659)

-8x - 20 = 25 1/3
-8x = 45 1/3 {added 20 to both sides}
-8x = 136/3 {changed 45 1/3 to improper fraction}
x = -17/3 {divided both sides by -8, multiplying by reciprocal -1/8}

For more help from me, visit: www.algebrahouse.com